OpenStudy (lgbasallote):

i was given this: \[f(t) = \cases{3 \quad t < 2 \\ 1 \quad 2

OpenStudy (lgbasallote):

look at all the beautiful latex lol

OpenStudy (lgbasallote):

the final answer is \[\Large \frac 3s - \frac{2e^{-2s}}{s} + e^{-5s} \left[ \frac{1}{s^2} + \frac 4s \right] + e^{-8s} \left[\frac{1}{5s^3} - \frac{1}{s^2} + \frac 8s \right]\]

OpenStudy (lgbasallote):

^just in case it cant be read earlier

OpenStudy (anonymous):

the first equation (step function) you got seems to be little bit wrong.

OpenStudy (lgbasallote):

the 3 - 2u thingy? that's right...i confirmed it with my notes

OpenStudy (anonymous):

yes 3-2u thing . let me check it again then.

OpenStudy (lgbasallote):

another thing i'd like to ask btw...do i write the final answer as y(s) = or y(t) =?

OpenStudy (anonymous):

it should in in the S domain since we are finding Laplace (from t -> s)

OpenStudy (lgbasallote):

so y(s)?

OpenStudy (anonymous):

F(s) i guess because f(t) is given !

OpenStudy (lgbasallote):

heh same shiz...anyway...

OpenStudy (lgbasallote):

why do you think my equation is wrong?

OpenStudy (anonymous):

i think it should be 3u(t-2)

OpenStudy (lgbasallote):

but the graph would cross the 0 right?

OpenStudy (lgbasallote):

 something like tha?